Erica age is 8 and her mother is 42. How many years time will her mother be three times as old as her daughter
Answer by stanbon(75887)

Erica age is 8 and her mother is 42. How many years time will her mother be three times as old as her daughter
------------------------
Let the answer to your problem be "x"
Equation:
42+x = 3(8+x)
42+x = 24+3x
2x = 18
x = 9
In 9 yrs Erica will be 8+9 = 17 ; her mother will be 42+9 = 51
51 is 3 times 17
---------------
Cheers,
Stan H.
----------------
-----------